3 blocks of equal mass 3kgs hang from a fixed pulley.mass A is on one side and B and C on another.Find the tension in the string connecting B and C and why?

Grade:12

4 Answers

Chetan Mandayam Nayakar
12 years ago

B and C accelerate downward at (2*3 -3)g/9 =g/3, let the tension in string connecting B and C be T.

3*g-T=3*(g/3)

solve for T

Ashwin Muralidharan IIT Madras
12 years ago

That gives the tension in the string connecting Block B and C.

 

Regards,

Ashwin (IIT Madras)

 

Pranjal Mittal
12 years ago

first we take a+b+c as a system

(6g-3g)/9=a

a=g/3

now 

force on C

3g-T=3*g/3

T=2g

neeraj agarwal
11 years ago

accln=g/3 =>3g-T=3g/3                                                                                                                                                                                         T=2g